OpenStudy (anonymous):

i promise a medal What is the rate of change between the interval of x = pi over 4 and x = pi over two

OpenStudy (anonymous):

2

OpenStudy (anonymous):

._.

OpenStudy (anonymous):

answer choices -8/pi -pi/8 -32/pi -pi/32

OpenStudy (anonymous):

rate of change of what?

OpenStudy (anonymous):

wait there is a graph

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

hold up lemme think

OpenStudy (anonymous):

f(b)−f(a)/b−a use this formula

OpenStudy (anonymous):

This is sine function with: amplitude = 4 frequency = 2π/3 vertical shift = −3 f(x) = 4 sin(3x) − 3 http://fooplot.com/#W3sidHlwZSI6MCwiZXEi... OK, after finding the actual function, using graph to find f(π/4) is a little misleading: f(π/4) = 4 sin(3π/4)−3 = 4(√2/2)−3 = 2√2−3 Rate of change = (f(π/2) − f(π/4)) / (π/2 − π/4) = (−7 − (2√2−3)) / (π/4) = (−4−2√2) / (π/4) = (−16−8√2) / π ≈ −8.694223444

OpenStudy (anonymous):

can u figure out the answer from this?

OpenStudy (anonymous):

-8/pi

OpenStudy (anonymous):

........

OpenStudy (anonymous):

ya correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

What is the logarithmic form of the equation e5x ≈ 4768?

OpenStudy (anonymous):

i have no idea XD

OpenStudy (anonymous):

@pooja195 @pooja195 @pooja195

OpenStudy (anonymous):